# A Fundamental Trade-off in Aligned Language Models and its Relation to Sampling Adaptors

### v1

@@ 5 Theoretical Results
We are now ready to discuss the theoretical contributions of this paper. In § 5.1 we begin with a formal argument that there exists a fundamental trade-off between the average log-probability under the prior and the average reward for corpora sampled from an aligned language model. Then, in § 5.2, we show how sampling adaptors, by shifting probability mass, can choose a point on this trade-off. In § 5.3, we conclude the section with a discussion of how this trade-off leads to an emergence of Simpson’s paradox.
@@ 5.1 A Fundamental Trade-off
$\Big{|}{\mathrm{H}}({\boldsymbol{Y}},{A}={\textsc{+}})+\frac{\log{q_{{\textsc{+}}}}({\mathcal{Y}})}{N}\Big{|}<\varepsilon\Big{\}}$
where ${\mathcal{Y}}$ is a corpus of strings and $\log{q_{{\textsc{+}}}}({\mathcal{Y}})=\sum_{n=1}^{N}\log{q_{{\textsc{+}}}}({{\boldsymbol{y}}^{(n)}})$ for some $\epsilon>0$. In words, ${T_{N}^{\varepsilon}}({q_{{\textsc{+}}}})$ is the set of corpora of size $N$, i.e., bags of strings, each sampled from ${q_{{\textsc{+}}}}$ with average information content $\frac{\log{q_{{\textsc{+}}}}({\mathcal{Y}})}{N}$ close to the pointwise joint entropy ${\mathrm{H}}({\boldsymbol{Y}},{A}={\textsc{+}})$.
@@
The full derivation can be found in § A.1. What Eq. 9 says is that if we observe a set of $N$ strings, the probability that the corpus lies outside ${T_{N}^{\varepsilon}}({q_{{\textsc{+}}}})\rightarrow 0$ as $N\rightarrow\infty$. Equivalently, this is to say that the sample entropy $-\frac{\log{q_{{\textsc{+}}}}({\mathcal{Y}})}{N}$ collapses around the entropy ${\mathrm{H}}({\boldsymbol{Y}},{A}={\textsc{+}})$ with high probability when $N$ is large.
The above derivation is standard. However, what is less standard is the application of Bayes’s rule to show that strings in the typical set display a fundamental trade-off.
@@
Prop. 1 follows relatively straightforwardly from Eq. 8 and Eq. 9, when one observes from Eq. 4 that ${q_{{\textsc{+}}}}({\boldsymbol{y}})\propto{p}({\boldsymbol{y}})\exp\left({{r}}({\boldsymbol{y}})\right)$, which implies $\log({q_{{\textsc{+}}}})=\log{p}({\boldsymbol{y}})+{{r}}({\boldsymbol{y}})+$ constant. Recall that corpora in the typical set ${T_{N}^{\varepsilon}}({q_{{\textsc{+}}}})$ have average information content $\frac{\log{q_{{\textsc{+}}}}({\boldsymbol{y}})}{N}$ close to the constant pointwise joint entropy ${\mathrm{H}}({\boldsymbol{Y}},{A}={\textsc{+}})$ (Eq. 8). That is, typical corpora by definition exhibit a trade-off between average log-probability under the prior $\frac{\log{p}({\boldsymbol{y}})}{N}$ and average reward $\frac{{{r}}({\boldsymbol{y}})}{N}$. Then, due to Chebyshev’s inequality in Eq. 9, we have $\Big{|}C+\frac{\log{p}({\mathcal{Y}})}{N}+\frac{{{r}}({\mathcal{Y}})}{\beta N}\Big{|}<\varepsilon$ with probability at least $\left(1-\frac{{\mathbb{V}}({I})}{N\varepsilon^{2}}\right)$ for all $N$ and $\varepsilon>0$. When $N\geq\frac{{\mathbb{V}}({I})}{\delta\varepsilon^{2}}$ for some $\delta={\mathcal{O}}(\frac{1}{N})$, the above holds with probability at least $(1-\delta)$, i.e., with high probability, and we arrive at the proposition. The full proof is provided in § A.2. ∎
@@
Prop. 1 relies on two key assumptions. First, that ${q_{{\textsc{+}}}}$ has finite entropy.777 In general, this is not true. See App. B for an example. Second, that the variance of information content of a string is also finite, i.e., ${\mathbb{V}}({I})<+\infty$. We argue that neither of these assumptions are limiting in practice because we show in Prop. 3 and Prop. 5 that they hold for all Transformer-based language models, which constitute the base architecture for most modern models (Brown et al., 2020; Touvron et al., 2023).
@@
We remark that there exists a tighter bound for the probability–quality trade-off than the ${\mathcal{O}}(\frac{1}{N})$ one in Eq. 9 for specific types of language models. Specifically, we show in App. C that for transformer and $n$-gram based language models, the probability that sampled corpora land in the typical set ${T_{N}^{\varepsilon}}({q_{{\textsc{+}}}})$ and exhibit the trade-off grows exponentially quickly, i.e., the bound is ${\mathcal{O}}\left(\exp(-cN)\right)$ for some constant $c\in{\mathbbm{R}_{+}}$.
@@ 5.2 Controlling the Trade-off
Ideally, we would like to choose how much probability we trade for quality when sampling corpora from an aligned model. After all, depending on the context, it may be desirable to extract higher-reward text (e.g., to improve alignment) or lower-reward text (e.g., to combat overfitting of the reward function; Azar et al., 2023; Gao et al., 2023; Wang et al., 2024; He et al., 2024).
We can leverage sampling adaptors to exercise this control. Sampling adaptors only modify string probability under the prior. That is, we can show:
${{\tilde{{p}}}_{+}}({\boldsymbol{y}})$
$=\frac{{\boldsymbol{\alpha}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}})\prod_{t=1}^{|{\boldsymbol{y}}|}{\boldsymbol{\alpha}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({{y}_{t}})}{{{Z}_{\boldsymbol{\alpha}}({\boldsymbol{y}})}}$
$\propto\left(\dots\right){p}({\textsc{eos}}\mid{\boldsymbol{y}})\prod_{t=1}^{|{\boldsymbol{y}}|}{p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}}){p}({A}={\textsc{+}}\mid{\boldsymbol{y}})$ (11)
where ${{\tilde{{p}}}_{+}}$ is the resultant distribution when applying a sampling adaptor ${\boldsymbol{\alpha}}$ to an aligned model ${{p}_{{\textsc{+}}}}$, and ()($\dots$)( … ) refers to a series of truncation functions applied to ${p}({{y}_{t}}\mid{{\boldsymbol{y}}_{<t}})$ at all $t$. See App. D for a full derivation.
§ 5.2 says that the effect of applying a sampling adaptor ${\boldsymbol{\alpha}}$ to ${{p}_{{\textsc{+}}}}$ is akin to applying ${\boldsymbol{\alpha}}$ to the prior language model ${p}$ and then multiplying the result by the likelihood that the generated string aligns with human preferences, i.e., the effects of the sampling adaptor can be pushed to the prior. This follows from Eq. 6, Bayes’ rule, and the fact that many sampling adaptors can be decomposed into a series of scaling and truncation operations.
Given the probability–quality trade-off, this implies that we can use sampling adaptors to control the average log-probability of sampled corpora, which then determines the average reward of generated text. For example, we could use temperature sampling with a high temperature to produce lower probability (and thus higher reward) strings.
Importantly, for a sampling adaptor to correctly control the trade-off, we require that the constant term $C\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}{\mathrm{H}}({\boldsymbol{Y}}\mid{A}={\textsc{+}})-\log{{Z}({\textsc{+}})}$ in Prop. 1 remains unchanged. Otherwise, the trade-off induced by ${{\tilde{{p}}}_{+}}$ would be different to that of ${{p}_{{\textsc{+}}}}$. Because the naive application of a sampling adaptor modifies the distribution to be sampled from, we require the IMHA correction introduced in § 4. With it, we can derive true samples of ${{p}_{{\textsc{+}}}}$ with a choice of sampling adaptor that controls the average log-probability of sampled corpora. As we demonstrate in § 6.2, using this procedure with different sampling adaptors allows us to choose how much log-probability we trade for reward when sampling from ${{p}_{{\textsc{+}}}}$.
@@ 6 Experimental Setup
We conduct two experiments in the empirical portion of this paper. First, in § 6.1 we validate the predictions of Prop. 1 with toy language and reward models. Then, in § 6.2 we demonstrate that this trade-off exists in practice for open-sourced RLHF-tuned models and that common sampling adaptors can control where on this trade-off the corpus of generated text will lie. We also examine models aligned with Direct Preference Optimization (DPO; Rafailov et al., 2023) in App. E, as RLHF and DPO have the same objective.
@@ 6.1 A Toy Experiment
The trade-off described in Prop. 1 fundamentally arises as a consequence of typicality and the fact that ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})\propto{p}({\boldsymbol{y}})\big{(}\exp{{r}}({\boldsymbol{y}})\big{)}$. We aim to demonstrate these theoretical principles with an easily reproducible toy experiment, where we model these distributions over a finite set of objects ${x}\in{D}$. That is, we will show the existence of the trade-off using the toy models ${{p}_{{\textsc{+}}}}({x})$, ${p}({x})$ and ${{r}}({x})$.
@@
We set $|{D}|=1,000$ and model ${{p}_{{\textsc{+}}}}({x})$ by sampling a distribution from the Dirichlet distribution. Then, we create the prior ${p}({x})$ by applying the softmax to a scaled and noised version of this distribution. That is, we define ${p}(\cdot)\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}\mathrm{softmax}\big{(}{{p}_{{\textsc{+}}}}(\cdot)^{\frac{1}{{\tau}}}+\epsilon)\big{)}$ where $\epsilon\sim U(-\kappa,\kappa)$ and ${\tau}\in{\mathbbm{R}_{+}}$, $\kappa\in{\mathbbm{R}_{+}}$ are our hyperparameters. We then define ${{r}}({x})$ analogously to Eq. 4 with ${{r}}({x})=\log\frac{{{p}_{{\textsc{+}}}}({x})}{{p}({x})}$. The distributions of ${{p}_{{\textsc{+}}}}({x})$, ${p}({x})$ and ${{r}}({x})$ over the domain ${D}$ can be seen in App. F. To induce Simpson’s paradox, we tune $\kappa$ and ${\tau}$ such that they are positively correlated, shown in Fig. 1 on the left.
@@ 6.2 The Trade-off in Practice
Here we demonstrate the existence of the probability–quality trade-off with an open-sourced aligned language model based on the Llama 2 family (Touvron et al., 2023). Using sampling adaptors we sample a corpus ${\mathcal{Y}}$ of 2,000 texts from an RLHF-tuned model ${q_{{\textsc{+}}}}$. Towards this, we randomly choose 1,000 prompts using the helpfulness dataset from Perez et al. (2022) and for each prompt, we produce two generations. Then, for every string in this corpus, we obtain its log-probability under the prior language model $\log{p}({\boldsymbol{y}})$ and its reward ${{r}}({\boldsymbol{y}})$. The prior and reward models are the same as those used to train ${q_{{\textsc{+}}}}$ in an RLHF scheme. We repeat this using five sampling adaptors at five temperatures, totaling 25 sampling schemes and thus $50,000$ $\big{(}\log{p}({\boldsymbol{y}}),{{r}}({\boldsymbol{y}})\big{)}$ pairs. To observe the trade-off, we compute the Pearson and Spearman’s correlation between $\log{p}({\boldsymbol{y}})$ and ${{r}}({\boldsymbol{y}})$ at the string level, and between $\frac{\log{p}({\mathcal{Y}})}{N}$ and $\frac{{{r}}({\mathcal{Y}})}{N}$ at the corpus level.
@@
To compute corpus-level correlations we require a lot of data points of the average log-probability and average reward. However, because sampling multiple corpora is prohibitively expensive, we use the IMHA with standard bootstrap resampling (Bradley Efron, 1994) to create multiple corpora for each of the 25 sampling schemes. Given a corpus of strings generated from ${q_{{\textsc{+}}}}$ with a sampling adaptor ${\boldsymbol{\alpha}}$, we resample uniformly with replacement $N$ times, accepting and rejecting each as described in Algorithm 1. This gives us a resampled corpus ${\mathcal{Y}}^{\prime}$. Then, we compute the average log-likelihood $\frac{\log{p}({\mathcal{Y}}^{\prime})}{N}$ and average reward $\frac{{{r}}({\mathcal{Y}}^{\prime})}{N}$. We do this 2,000 times per sampling scheme, giving us a total of 50,000 $\left(\frac{\log{p}({\mathcal{Y}})}{N},\frac{{{r}}({\mathcal{Y}}^{\prime})}{N}\right)$ pairs, which we then use to we compute the corpus-level correlations. We set $N=200,000$ as preliminary experiments showed that for $N\geq 200,000$ the IMHA converges, i.e, the autocorrelation measured with Cramer’s V falls to $<0.10$.
Sampling Adaptors.
The five sampling adaptors we examine are: top-${k}$ sampling (Fan et al., 2018) for ${k}\in{\{30,50\}}$, nucleus sampling (Holtzman et al., 2020) for ${\pi}\in{\{0.90,0.95\}}$, $\eta$ sampling (Hewitt et al., 2022) and locally-typical sampling (Meister et al., 2023b). For each setting, we examine five temperatures ${\tau}\in{\{0.5,0.75,1.0,1.25,1.5\}}$. This gives us a total of 25 settings that cover various real-world use cases. As a baseline, we also include ancestral sampling.

### v2

@@ 5 Theoretical Results
We are now ready to discuss the theoretical contributions of this paper. In § 5.1 we begin with a formal argument that there exists a fundamental trade-off between the average log-probability under the prior and the average reward for corpora sampled from an aligned language model. Then, in § 5.2, we show how sampling adaptors, by shifting probability mass, we can choose any point on this trade-off. In § 5.3, we conclude the section with a discussion of how this trade-off leads to an emergence of Simpson’s paradox.
@@ 5.1 A Fundamental Trade-off
$\Big{|}{\mathrm{H}}({\boldsymbol{Y}},{A}={\textsc{+}})+\frac{\log{q_{{\textsc{+}}}}({\mathcal{Y}})}{N}\Big{|}<\varepsilon\Big{\}},$
where ${\mathcal{Y}}$ is a corpus of strings and $\log{q_{{\textsc{+}}}}({\mathcal{Y}})=\sum_{n=1}^{N}\log{q_{{\textsc{+}}}}({{\boldsymbol{y}}^{(n)}})$ for some $\varepsilon>0$. In words, ${T_{N}^{\varepsilon}}({q_{{\textsc{+}}}})$ is the set of corpora of size $N$, i.e., bags of strings, each sampled from ${q_{{\textsc{+}}}}$ with average information content $\frac{\log{q_{{\textsc{+}}}}({\mathcal{Y}})}{N}$ close to the pointwise joint entropy ${\mathrm{H}}({\boldsymbol{Y}},{A}={\textsc{+}})$.
@@
The full derivation can be found in § B.1. What Eq. 11 says is that if we observe a set of $N$ strings, the probability that the corpus lies outside ${T_{N}^{\varepsilon}}({q_{{\textsc{+}}}})\rightarrow 0$ as $N\rightarrow\infty$. Equivalently, this is to say that the sample entropy $-\frac{\log{q_{{\textsc{+}}}}({\mathcal{Y}})}{N}$ collapses around the entropy ${\mathrm{H}}({\boldsymbol{Y}},{A}={\textsc{+}})$ with high probability when $N$ is large.
The above derivation is standard. However, what is less standard is the application of Bayes’ rule to show that strings in the typical set display a fundamental trade-off.
@@
author=ryan,color=violet!40,size=,fancyline,caption=,]need to check the rest of this proof Prop. 1 follows relatively straightforwardly from Eq. 10 and Eq. 11, when one observes from Eq. 4 that ${q_{{\textsc{+}}}}({\boldsymbol{y}})\propto{p}({\boldsymbol{y}})\exp\left(\frac{1}{\beta}{{r}}({\boldsymbol{y}})\right)$, which implies $\log({q_{{\textsc{+}}}})=\log{p}({\boldsymbol{y}})+\frac{1}{\beta}{{r}}({\boldsymbol{y}})+\mathrm{constant}$. Recall that corpora in the typical set ${T_{N}^{\varepsilon}}({q_{{\textsc{+}}}})$ have average information content $\frac{\log{q_{{\textsc{+}}}}({\boldsymbol{y}})}{N}$ close to the constant pointwise joint entropy ${\mathrm{H}}({\boldsymbol{Y}},{A}={\textsc{+}})$ (Eq. 10). That is, typical corpora by definition exhibit a trade-off between average log-probability under the prior $\frac{\log{p}({\boldsymbol{y}})}{N}$ and average reward $\frac{{{r}}({\boldsymbol{y}})}{N}$. Then, due to Chebyshev’s inequality in Eq. 11, we have $\Big{|}C+\frac{\log{p}({\mathcal{Y}})}{N}+\frac{{{r}}({\mathcal{Y}})}{\beta N}\Big{|}<\varepsilon$ with probability at least $\left(1-\frac{{\mathbb{V}}({I})}{N\varepsilon^{2}}\right)$ for all $N$ and $\varepsilon>0$. When $N\geq\frac{{\mathbb{V}}({I})}{\delta\varepsilon^{2}}$ for some $\delta={\mathcal{O}}(\frac{1}{N})$, the above holds with probability at least $(1-\delta)$, i.e., with high probability, and we arrive at the proposition. The full proof is provided in § B.2. ∎
@@
Prop. 1 relies on two key assumptions. First, that ${q_{{\textsc{+}}}}$ has finite entropy.666 In general, this is not true. See App. C for an example. Second, that the variance of information content of a string is also finite, i.e., ${\mathbb{V}}({I})<+\infty$. We argue that neither of these assumptions are limiting in practice because we show in Prop. 3 and Prop. 5 that they hold for all Transformer-based language models, which constitute the base architecture for most modern models (Brown et al., 2020; Touvron et al., 2023).
@@
We remark that there exists a tighter bound for the probability–quality trade-off than the ${\mathcal{O}}(\frac{1}{N})$ one in Eq. 11 for specific types of language models. Specifically, we show in App. D that for transformer and $n$-gram based language models, the probability that sampled corpora land in the typical set ${T_{N}^{\varepsilon}}({q_{{\textsc{+}}}})$ and exhibit the trade-off grows exponentially quickly, i.e., the bound is ${\mathcal{O}}\left(\exp(-cN)\right)$ for some constant $c\in{\mathbbm{R}_{>0}}$.
@@ 5.2 Controlling the Trade-off
Ideally, we would like to choose how much probability we trade for quality when sampling corpora from an aligned language model. After all, depending on the context, it may be desirable to extract higher-reward text (e.g., to improve alignment) or lower-reward text (e.g., to combat overfitting of the reward function; Azar et al., 2023; Gao et al., 2023; Wang et al., 2024; He et al., 2024).
We can leverage globally normalized sampling adaptors (§ 4) to exercise this control. Globally normalized sampling adaptors directly select strings based on their string probability under the prior. That is, we can show:
${{\widetilde{{p}}}_{+}}({\boldsymbol{y}})$
$\propto{\boldsymbol{\gamma}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({\textsc{eos}}){\prod^{|{\boldsymbol{y}}|}_{t=1}}{\boldsymbol{\gamma}}\big{(}{{p}_{{\textsc{+}}}}(\cdot\mid{\boldsymbol{y}})\big{)}({{y}_{t}})$
$\propto\ \underbrace{\left(\dots\right){\prod^{|{\boldsymbol{y}}|}_{t=1}}\left(\dots\right)}_{\text{Functions of the Prior, ${p}$}}\,{{p}_{{\textsc{+}}}}({\boldsymbol{y}})$ (13)
where ${{\widetilde{{p}}}_{+}}$ is the induced distribution when applying a transform function ${\boldsymbol{\gamma}}$ to an aligned model ${{p}_{{\textsc{+}}}}$, and ()($\dots$)( … ) refers to the series of scaling and truncation functions that take the prior as input. See App. E for a full derivation.
§ 5.2 says that the effect of applying a globally normalized sampling adaptor to ${{p}_{{\textsc{+}}}}$ is akin to applying it to the prior language model ${p}$ and then multiplying the result by the likelihood that the generated string aligns with human preferences, i.e., the effects of the sampling adaptor can be pushed to the prior. This follows from Eq. 6, Bayes’ rule, and the fact that many transform functions used in globally and locally normalized sampling adaptors can be decomposed into a series of scaling and truncation operations.
Given the probability–quality trade-off, this implies that an appropriate choice of globally normalized sampling adaptor can be used to control the average log-probability of sampled corpora, which then determines the average reward of generated text. For example, we could use the globally normalized equivalent of temperature sampling with a high temperature to produce lower probability (and thus higher reward) strings. We demonstrate this control in § 6.2.
@@ 6 Experimental Setup
We conduct two experiments in the empirical portion of this paper. First, in § 6.1 we validate the predictions of Prop. 1 with toy language and reward models. Then, in § 6.2 we demonstrate that this trade-off exists in practice for open-sourced RLHF-tuned models and that globally normalized sampling adaptors can control where on this trade-off the corpus of generated text will lie. We also examine models aligned with Direct Preference Optimization (DPO; Rafailov et al., 2023) in App. F, as RLHF and DPO have the same objective.
@@ 6.1 A Toy Experiment
The trade-off described in Prop. 1 fundamentally arises as a consequence of typicality and the fact that ${{p}_{{\textsc{+}}}}({\boldsymbol{y}})\propto{p}({\boldsymbol{y}})\exp{{r}}({\boldsymbol{y}})$. We aim to demonstrate these theoretical principles with an easily reproducible toy experiment, where we model these distributions over a finite set of objects ${x}\in{D}$. That is, we will show the existence of the trade-off using the toy models ${{p}_{{\textsc{+}}}}({x})$, ${p}({x})$ and ${{r}}({x})$.
@@
We set $|{D}|=1,000$ and model ${{p}_{{\textsc{+}}}}({x})$ by sampling a distribution from the Dirichlet distribution. Then, we create the prior ${p}({x})$ by applying the softmax to a scaled and noised version of this distribution. That is, we define ${p}(\cdot)\mathrel{{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}}\mathrm{softmax}\big{(}{{p}_{{\textsc{+}}}}(\cdot)^{\frac{1}{{\tau}}}+\epsilon)\big{)}$ where $\epsilon\sim U(-\kappa,\kappa)$ and ${\tau}\in{\mathbbm{R}_{>0}}$, $\kappa\in{\mathbbm{R}_{>0}}$ are our hyperparameters. We then define ${{r}}({x})$ analogously to Eq. 4 with ${{r}}({x})=\log\frac{{{p}_{{\textsc{+}}}}({x})}{{p}({x})}$. The distributions of ${{p}_{{\textsc{+}}}}({x})$, ${p}({x})$ and ${{r}}({x})$ over the domain ${D}$ can be seen in App. G. To induce Simpson’s paradox, we tune $\kappa$ and ${\tau}$ such that they are positively correlated, shown in Fig. 1 on the left.
@@ 6.2 The Trade-off in Practice
Here we demonstrate the existence of the probability–quality trade-off with an open-sourced aligned language model based on the Llama 2 family (Touvron et al., 2023). Using locally normalized sampling adaptors, we sample a corpus ${\mathcal{Y}}$ of 2,000 texts from an RLHF-tuned model ${q_{{\textsc{+}}}}$. Towards this, we randomly choose 1,000 prompts using the helpfulness dataset from Perez et al. (2022) and for each prompt, we produce two generations. Then, for every string in this corpus, we obtain its log-probability under the prior language model $\log{p}({\boldsymbol{y}})$ and its reward ${{r}}({\boldsymbol{y}})$. The prior and reward models are the same as those used to train ${q_{{\textsc{+}}}}$ in an RLHF scheme. We repeat this using five locally normalized sampling adaptors at five temperatures, totaling 25 sampling schemes and thus $50,000$ $\big{(}\log{p}({\boldsymbol{y}}),{{r}}({\boldsymbol{y}})\big{)}$ pairs. To observe the trade-off, we compute the Pearson and Spearman’s correlation between $\log{p}({\boldsymbol{y}})$ and ${{r}}({\boldsymbol{y}})$ at the string level, and between $\frac{\log{p}({\mathcal{Y}})}{N}$ and $\frac{{{r}}({\mathcal{Y}})}{N}$ at the corpus level.
@@
To compute corpus-level correlations we require a lot of data points of the average log-probability and average reward. However, because sampling multiple corpora is prohibitively expensive, we use standard bootstrap resampling (Efron and Tibshirani, 1994) to create multiple corpora for each of the 25 sampling schemes. Given a corpus of strings generated from ${q_{{\textsc{+}}}}$ with a locally normalized sampling adaptor defined by the transform function ${\boldsymbol{\gamma}}$, we resample uniformly with replacement $N$ times, accepting and rejecting each as described in Algorithm 1. This gives us a resampled corpus ${\mathcal{Y}}^{\prime}$. Then, we compute the average log-likelihood $\frac{\log{p}({\mathcal{Y}}^{\prime})}{N}$ and average reward $\frac{{{r}}({\mathcal{Y}}^{\prime})}{N}$. We do this 2,000 times per sampling scheme, giving us a total of 50,000 $\left(\frac{\log{p}({\mathcal{Y}})}{N},\frac{{{r}}({\mathcal{Y}}^{\prime})}{N}\right)$ pairs, which we then use to we compute the corpus-level correlations. We set $N=200,000$ as preliminary experiments showed that for $N\geq 200,000$ the IMHA converges, i.e, the autocorrelation measured with Cramér’s V falls to $<0.10$.
Sampling adaptors
The five sampling schemes we examine are: top-${k}$ sampling (Fan et al., 2018) for ${k}\in{\{30,50\}}$, nucleus sampling (Holtzman et al., 2020) for ${\pi}\in{\{0.90,0.95\}}$, $\eta$ sampling (Hewitt et al., 2022) and locally-typical sampling (Meister et al., 2023b). For each setting, we examine five temperatures ${\tau}\in{\{0.5,0.75,1.0,1.25,1.5\}}$. This gives us a total of 25 settings that cover various real-world use cases. As a baseline, we also include ancestral sampling.